# On the Surprising Effectiveness of Spectral Clipping in Learning Stable Linear Dynamics

### v4

@@ 4 Empirical Analyses
In this section, we evaluate SC in terms of computation efficiency, stability, and expressivity when learning LDS from data and compare its performance against existing methods. Note that for all experiments, we clip the spectrum to $1$ (see Appendix. E for the effects of different $\varepsilon$ values).
@@ 4.1 Experimental Design
Computation Platform:
We conducted all experiments on Alienware Aurora R13, with 10-core 12th Gen Intel Core i7-12700KF 1.0-GHz CPU and a 32G RAM.
Baselines:
i) LS: vanilla least squares, ii) (CG): constraint generation [2], iii) WLS: weighted least squares [24], and iv) SOC: a characterization approach [26, 27].
Datasets:
We evaluate all approaches on the following datasets.
UCSD [1]: A dataset of 254 highway traffic videos. Each video has 48-52 frames and the size of each frame is $48\times 48$. To lower the dimension, we apply the SVD method on each sequence to obtain a subspace dimensions $r\in\{3,30\}$, as similarly done in [26].
DTDB [34]: A dataset of 285 dynamic texture movements, such as turbulence, wavy motions, etc. Each video consists of varying frames, and we also apply the SVD method to lower the state dimension to $r=300$.
Metrics:
We evaluate all approaches in terms of: i) Computation time: Time taken to train a stable LDS, ii) Memory Usage: Memory used during the entire computation process, and iii) Reconstruction error: The $L1$ distance between generated state sequences and ground truth states.
@@ 4.2 SC significantly reduces computation & memory burden
In this subsection, we compare the training efficiency among each approach. We first report the average computation time across the 254 input sequences of the UCSD dataset at each dimension in Fig. 1 (a). The results show that SC is significantly more time-efficient than WLS, CG, and SOC. This is because that SC only needs the eigen-decomposition and eigen-recomposition to analytically compute the stable LDSs, while others require to iteratively find a solution. Note that though LS is the most time-efficient, it can not provide with stability.
In Fig. 1 (b), we then report the results on the DTDB dataset. The results also indicate that the computation time of SC is significantly and consistently shorter than that of SOC, showing that SC is more scalable w.r.t. the input dimensions. Similarly, the memory usage results reported in Fig. 1 (c) and Fig. 1 (d) suggest that SC requires significantly less memory compared to CG, WLS, and SOC.
@@
LS CG [2] WLS [24] SOC [26, 27] SC
Time complexity -- $O(n^{6})$ $O(n^{6})$ $O(n^{3})$ $O(n^{3})$
Space complexity $O(n^{2})$ $O(n^{4})$ $O(n^{4})$ $O(n^{2})$ $O(n^{2})$
Table 1: Space and Time complexity of each method. $n$ is the system dimension. For time complexity, since all methods modify the original LS solution to enforce stability, we compare only the cost incurred after solving LS. For iterative methods (CG, WLS, and SOC), the shown time complexity corresponds to a single iteration.
In addition to empirical results on the two datasets, we report the time and space complexity of each method in Table. 1, further supporting the observation that SC is as or more efficient in both time and memory than the other methods. For time complexity, since all methods modify the original LS solution to enforce stability, we compare only the cost incurred after solving LS. Note that CG, WLS, and SOC are iterative algorithms, and the table reports the time complexity per iteration. In contrast, SC is non-iterative, and its overall time complexity is shown.
4.3 SC effectively identifies the underlying LDSs
In this subsection, we compare the prediction accuracy of the identified LDSs by each approach. We first present the average reconstruction error across all states at each time step for the UCSD dataset in Fig. 2 (left). For illustrative purposes, we select $r=3,20$ and $30$ as examples. It is clearly shown that as the input dimension increases (e.g., $r=30$), LS fails to reconstruct the input states, with the error increasing significantly over time. A similar trend can be observed in Fig. 3, where we show the average reconstruction error over all states and time steps of each method at each training dimension. In both cases, SC consistently achieves results comparable to CG, WLS, and SOC, while being orders of magnitude faster for computation and highly efficient in memory usage (see Section. 4.2).
We then show the average reconstruction error for the DTDB dataset in Fig. 2 (right). Note that due to the high dimension of the state space in this dataset, only LS, SOC, and SC are applicable. Both CG and WLS involve solving a Kronecker product, which has a significantly large space complexity (i.e., $O(n^{4})$ shown in Table. 1, $n$ is the size of the state space), leading to memory overflow.
Given the varying time steps of the input sequences, the reconstruction error is averaged over the time percentiles. The results indicate that SC outperforms SOC since it achieves a lower reconstruction error. Note that though the LS approach achieves the lowest error within the training data horizon, it does not guarantee stability, which is crucial for long-term prediction (see Section. 4.4).
@@ 4.3 SC effectively identifies the underlying LDSs
4.4 SC enables stable long-term predictions
In this subsection, we demonstrate that the SC approach learns LDSs capable of making long-term dynamic predictions that extend beyond the training data horizon. To verify this, we extend the rollout horizon to $T=3000$ time steps for the 285 LDSs learned from the DTDB dataset. In this case, an accurate LDS should be able to repeat the dynamical pattern learned from the training data without either blowing up or converging during the extended rollout horizon.
We first select two example sequences (i.e., flushing toilet and running stream) and show the qualitative results in Fig. 6. These results indicate that LS approach mostly learns the LDSs that blow up during the extended horizon, as it does not provide with stability guarantee. Conversely, the SOC approach tends to learn LDSs that converge to zeros, indicating a lack of dynamic behavior. In Table. 2, we report the spectral radius of the learned system matrix for each method. The results support the conclusion that LS leads to blow-ups due to a spectral radius greater than one, while SOC results in vanishing values as its spectral radius remains strictly less than one.
LS SOC [26, 27] SC
DTDB Sequence 185 1.0025 0.9999 1.0000
DTDB Sequence 205 1.0022 0.9999 1.0000
Table 2: Spectrum radii of the system matrices learned with each method.
Figure 4: DTDB Sequence 185: A flushing toilet.
Figure 5: DTDB Sequence 205: A running stream.
Figure 6: Qualitative results for 2 example sequences from the DTDB dataset.
We then define several quantitative evaluation metrics, which are divided into two categories.
Quality (NIQE score [44]): A commonly used image quality evaluation metric. This metrics helps determine if the predictions of the LDSs are still meaningful. Note that images with extremely low quality (for example, when most pixel values converge to 0 or blow up) can produce unbounded NIQE scores. We substitute those unbounded scores with $10^{5}$ to provide a bounded statistical summaries.
Dynamics (Moving Ratio  [^1] Moving is defined as $||F_{M}-F_{M-2}||_{1}>T$, where $\{F_{i}\}_{i=1}^{M}$ is a sequence of $M$ image frames and $T=9$.): The ratio of sequences in which the final frames still exhibit the dynamic pattern This characterization aims to determine if the system evolution of the LDSs still reflects the dynamic pattern.
@@ 4.4 SC enables stable long-term predictions
SOC [26, 27] 832.8063 4243.4279 22.11%
SC 18.8769 1075.0140 54.39%
[…]
Table 3: Quantitative metrics evaluating the predictions of each method on the DTDB dataset.
In Table. 3, we report the quantitative ratios in terms of the defined metrics over the 285 sequences in the DTDB dataset. Note that the values for LS method are marked in gray as this method does not guarantee stability. The results indicate that SC best preserves long-term dynamic patterns while maintaining high prediction accuracy, as evidenced by its superior image quality—particularly at 500-step predictions, which are nearly comparable to the training set images.
@@ 4.5 SC’s benefits extend to nonlinear systems when combined with the Koopman Operator
In this subsection, we demonstrate that when combined with the Koopman Operator theory [7], SC can also be effectively and efficiently applied to highly nonlinear robot manipulation tasks. For this, we introduce a highly nonlinear robot manipulation dataset.
DexManip [13, 45]: A dataset of Adroit Hand performing two dexterous manipulation tasks in MUJOCO simulation environment [46]. Though the original state evolution is nonlinear, but as shown in [13], via lifted to a higher-dimension Hilbert space (e.g., $r\approx 750$), the system can be considered linear in the lifted state space. Two tasks are briefly described as follows: i) Tool Use: The robot hand is tasked with picking up the hammer and driving the nail into the board placed at a random height, and ii) Object Relocation: The robot hand is tasked with moving a cylinder to a random target location (see Appendix. B for the details of each dataset). Similar to the DTDB dataset, only LS, SOC, and SC are applicable due to the high dimensionality of the lifting space.
Computation Time (sec) Memory Usage (MB) Task Success Rate (%) Safety Rate (%)
@@
We first report the computation time and memory usage in Table. 4. As seen, SC also alleviates the computational and memory burden compared to the SOC method. Note that the values for LS method are marked in gray as this method does not guarantee stability. We then show the average reconstruction error on the DexManip datasets in Fig. 7 (a) & (b), where SC demonstrates the best performance and LS results in significantly larger errors on the Tool dataset due to its lack of stability.
Figure 7: (a): Average reconstruction error on Tool dataset; (b): Average reconstruction error on Reloc dataset; (c): Average reconstruction error on Tool_short dataset; (d): Average reconstruction error on Reloc_with_failures dataset. Solid lines represent mean values, and shaded areas indicate quartile values over the rollout trajectories.
To further underscore the importance of ensuring stability when learning LDSs for robotic dexterous manipulation tasks in the DexManip datasets, we also compute the task success rates and the safety rates (both reported in Table. 4) for each task when tracking the generated robot hand trajectories. The definition of these two metrics are included in Appendix. B.
@@
4.6 SC can learn stable LDSs from shorter and corrupted datasets
In real-world settings, the collected demonstrations may be imperfect, unlike those in the original DexManip datasets. Hence, we introduce two additional datasets: (i) Tool_short, where each demonstration is shorter, posing challenges in learning stable LDSs with a limited horizon, and (ii) Reloc_with_failures, which includes a few failed demonstrations (see Appendix. B for more details).
Computation Time (sec) Memory Usage (MB) Task Success Rate (%) Safety Rate (%)
@@ 4.6 SC can learn stable LDSs from shorter and corrupted datasets
Table 5: SC can efficiently learn stable manipulation skills from shorter and corrupted datasets.
@@
In Table. 5, we report the computational time and memory usage, and it is evident that SC is more efficient than SOC. In Fig. 7 (c) & (d), we show the average reconstruction error over the original uncorrupted datasets. From these results, we can observe that the LDSs learned by the LS approach tend to be highly unstable, as evidenced by the significantly increasing error over time. This indicates the importance of enforcing stability in learned LDSs to effectively handle corrupted demonstrations.
A similar conclusion can be drawn from the task success rates and safety rates (both reported in Table. 5), where SC demonstrates significantly superior performance.
@@
We also reveal the connections between SC and the Koopman Operator’s spectral properties. Following the algorithm described in [33], we can rollout the systems with only subsets of the Koopman modes and visualize the hand motions to investigate the effectiveness of SC upon each subset.
Specifically, for each eigenvalue $\lambda_{i}\ (i=1,...,n)$ of Koopman matrix $K$, we can find corresponding eigenvector $\mathbf{v_{i}}$ (i.e., the mode) and eigenvector $\mathbf{w_{i}}$ of adjoint matrix $K^{*}$, such that $K\mathbf{v_{i}}=\lambda_{i}\mathbf{v_{i}}\text{, and }K^{*}\mathbf{w_{i}}=\overline{\lambda_{i}}\mathbf{w_{i}}$. Then following [33], we construct the corresponding eigenfunction $\varphi_{i}:\mathbb{R}^{n}\rightarrow C$, defined as $\varphi_{i}(\mathbf{z})=\langle\mathbf{z},\mathbf{w}_{i}\rangle$. By doing so, the rollout of eigenfunctions can be represented by $\varphi_{i}(K\mathbf{z})=\lambda_{i}\varphi_{i}(\mathbf{z})$, and the states can be reconstructed by $\mathbf{z_{k}}=\sum_{i=1}^{n}\lambda_{i}^{k}\varphi_{i}(\mathbf{z_{0}})\mathbf{v_{i}}$. Based on this formula, specific modes can be selected for rollouts.
In our experiments, we first computed the Koopman matrix $K_{f}$ from $Reloc\_with\_failures$ dataset. We then sorted the eigenvalues by their norms from largest to smallest and divided them into two subsets for rollouts: (i) Unstable Set: the first 52 Koopman modes with unstable eigenvalues, which have norms greater than 1, and (ii) Stable Set: the remaining 707 modes. We then visualized the robot hand motions of each subset via the eigenfunction rollouts in Fig. 8, respectively (see the supplementary video for full policy rollouts). It can be seen that the rollout of the Unstable Set before clipping clearly generate unstable hand motions (see upper line of the green box), which results in the low success rate and safety rate when combined with Stable Set for the final task performance (as shown in Table. 5). Further, we clipped the matrix $K_{f}$ and visualized the rollouts of the clipped Unstable Set in Fig. 8. It is obvious that after clipping, the robot hand motions become stable (see bottom line of the green box), enabling the composite system to safely complete the task (as also shown in Table 5). The full policy rollout is also available in the supplementary video.
4.7 SC can identify systems with control inputs
In this subsection, we demonstrate the comparable capability of SC in identifying stable LDSs with control inputs (see Appendix. C for details). To this end, we first introduce another dataset that includes the control commands.
Franka Panda [35]: A dataset of Franka Panda Robot Arm’s movements. The robot state space has 17 dimensions ($r=17$), including the coordinates of end effector ($\mathbb{R}^{3}$), joint angle values ($\mathbb{R}^{7}$), and joint angular velocities ($\mathbb{R}^{7}$), and the control input has $\mathbb{R}^{7}$ joint torques. This dataset consists of 3100 time steps of control inputs (randomly generated) and the corresponding robot states.
We first evaluate the long-term prediction accuracy of each approach. For this, we first command the random control inputs in simulation with ground-truth system dynamics to generate test trajectories. We then rollout the trained $A$ and $B$ matrices with the same control inputs, and then compute the average $L2$ distance between the predicted and ground-truth end-effector positions for each method. From the results shown in Table. 6, We observe that while SC performs worse than CG and SOC when $d=100$, its performance becomes comparable as $d$ increases. Notably, SC achieves this with significantly lower computation time and memory usage.
Additionally, we use the trained $A$ and $B$ matrices to construct an LQR controller, and use such controller to track a figure-8 pattern in simulation (similar in [26]). How accurately the figure is tracked indicates how well the learned LDS aligns with the actual dynamical system with control inputs. For each approach, we repeat this process five times with different initial positions of the end effector. We show the qualitative tracking performances in Appendix. D, where SC demonstrates performance comparable to that of other methods.
LS CG [2] WLS [24] SOC [26, 27] SC
$d=100$ $0.0748\pm 0.1218$ $0.0166\pm 0.0066$ $0.3390\pm 0.2059$ $0.0225\pm 0.0131$ $0.0693\pm 0.1370$
$d=2000$ $0.0207\pm 0.0173$ $0.0154\pm 0.0058$ $0.3051\pm 0.1853$ $0.0154\pm 0.0058$ $0.0183\pm 0.0138$
$d=3000$ $0.0164\pm 0.0083$ $0.0155\pm 0.0058$ $0.2845\pm 0.1501$ $0.0155\pm 0.0058$ $0.0156\pm 0.0057$
5 Limitations and Future Work
Though spectrum clipping (SC) demonstrates its efficacy on identifying the stable LDSs over various datasets, there are still opportunities for future study. First, the performance on the dataset with control inputs is not yet fully comparable to other state-of-the-art approaches, indicating potential for algorithmic improvements. Second, while we empirically demonstrate the effectiveness of SC, a deeper theoretical analysis will likely offer valuable insights.
6 Conclusion
In this paper, we demonstrate that Spectrum Clipping is an efficient data-driven method for identifying stable LDSs. Through extensive experiments across multiple datasets, we show that it is not only significantly more time- and memory-efficient but also achieve superior performance. Furthermore, we demonstrate that it also excels in long-term dynamics prediction. Additionally, it can be applied to the highly nonlinear robot manipulation tasks with the Koopman Operator theory and the linear dynamical systems with control inputs.

### v6

@@ 5 Empirical Analyses
We evaluated SC in terms of computation efficiency, stability, and expressivity when learning LDS from data and compare its performance against existing methods. Note that for all experiments, we use $\epsilon=0$ unless indicated otherwise (see Appendix. C for the effects of different $\varepsilon$ values).
@@ 5.1 Experimental Design
Compute: We conducted all experiments on Alienware Aurora R13, with 10-core 12th Gen Intel Core i7-12700KF 1.0-GHz CPU with 32G RAM.
Baselines: We compared SC against the following baselines: i) LS: vanilla least squares, ii) (CG): constraint generation (Boots et al., 2007), iii) WLS: weighted least squares (Huang et al., 2016), and iv) SOC: a characterization approach (Mamakoukas et al., 2020, 2023).
Datasets: We evaluated all approaches on the following datasets.
i) UCSD (Chan and Vasconcelos, 2005): This dataset contains 254 highway traffic videos, each consisting of 48–52 frames of size $48\times 48$. Following Mamakoukas et al. (2020). we apply SVD to each sequence and retain subspaces of dimension $r\in\{3,30\}$ and reduce dimensionality.
ii) DTDB (Hadji and Wildes, 2018): A dataset of 285 dynamic texture movements, such as turbulence, wavy motions, etc. Each video consists of varying number of frames. As we did for the UCSD dataset, we leveraged SVD to lower the state dimension to $r=300$.
iii) DexManip (Han et al., 2023; Rajeswaran et al., 2018b): A robotic dexterous manipulation dataset containing demonstrations of a 23-DoF multi-fingered Adroit hand performing four complex tasks in MUJOCO (Todorov et al., 2012).
iv) Franka Panda (Gaz et al., 2019): A dataset of Franka Panda Robot Arm’s movements as a consequence of different control inputs. Results are reported in Appendix E.
@@ 5.2 SC significantly reduces computation & memory burden
LS CG WLS SOC SC
Time complexity - $O(n^{6})$ $O(n^{6})$ $O(n^{3})$ $O(n^{3})$
Space complexity $O(n^{2})$ $O(n^{4})$ $O(n^{4})$ $O(n^{2})$ $O(n^{2})$
Table 1: Space and Time complexity of each method with $n$ being the system dimension.
We begin by comparing the computational efficiency of each approach. We report the average computation time across the 254 input sequences of the UCSD dataset at each dimension in Fig. 2 (a). The results show that SC is significantly more time-efficient than WLS, CG, and SOC. In Fig. 2 (b), we report the results on the DTDB dataset. These results also indicate that the computation time of SC is significantly and consistently shorter than that of SOC, showing that SC is more scalable w.r.t. the input dimensions. Similarly, the memory usage results reported in Fig. 2 (c) and Fig. 2 (d) suggest that SC requires significantly less memory compared to CG, WLS, and SOC. The improved efficiency is a result of SC only requiring eigen decomposition and simple algebraic calculations, while others rely on iterative optimization. Though LS is the naturally the most time-efficient, it does not guarantee stability.
@@
In addition to empirical results on the two datasets, we report the time and space complexity of each method in Table. 1. For time complexity, since all methods modify the original LS solution to enforce stability, we compare only the cost incurred after solving LS. For iterative methods (CG, WLS, and SOC), the shown time complexity corresponds to a single iteration. In contrast, SC is non-iterative, and its overall time complexity is shown. These results further suggest that SC’s time and memory efficiency is comparable or better than those of the baselines.
5.3 SC does not sacrifice predictive accuracy
To evaluate if SC’s improved efficiency comes at the expense of predictive accuracy, we compared the prediction accuracy of LDSs produced by each approach. We first present the average reconstruction error across all states at each time step for the UCSD dataset in Fig. 4 (left). For illustrative purposes, we select $r=3,20$ and $30$ as examples. We find that as the input dimension increases (e.g., $r=30$), LS fails to reconstruct the input states, with the error increasing significantly over time. A similar trend can be observed in Fig. 3, where we show the average reconstruction error over all states and time steps of each method at each training dimension. In both cases, SC consistently achieves results comparable to CG, WLS, and SOC, while being orders of magnitude faster for computation and highly efficient in memory usage (as discussed in Section. 5.2).
In Fig. 4 (right), we report how prediction error varies across time for the DTDB dataset as length of the input sequence given to the model before predictions. Note that due to the high dimension of the state space in this dataset, only LS, SOC, and SC are applicable. Both CG and WLS involve solving a Kronecker product, which has a significantly large space complexity leading to memory overflow (i.e., $O(n^{4})$ as shown in Table. 1, $n$ is the size of the state space).
The results indicate that SC generates marginally more accurate predictions than SOC, and its error grows slower as the prediction horizon increases. While LS achieves the lowest error, it does not guarantee stability, a crucial property for reliable long-horizon predictions (see Section. 5.4).
@@ 5.3 SC does not sacrifice predictive accuracy
5.4 SC enables consistent long-horizon predictions
We evaluated that SC had the ability to make sensible long-horizon predictions that extend beyond the time-horizon of the training data. To verify this, we extend the rollout horizon to both $H=500$ and $H=3000$ time steps for the 285 LDSs learned from the DTDB dataset. Ideally, the learned system would be able to repeat the dynamical pattern learned from data, avoid both instability and convergence (black/static pixels) despite the extended rollout horizon.
Avg. NIQE ($H=500$) ($\downarrow$) Avg. NIQE ($H=3000$) ($\downarrow$) Moving Ratio ($\uparrow$)
@@ 5.4 SC enables consistent long-horizon predictions
SOC 832.8063 4243.4279 22.11%
SC ($\varepsilon=0$) 18.8769 1075.0140 54.39%
SC($\varepsilon=10^{-5}$) 18.8769 1075.0139 54.04%
SC($\varepsilon=10^{-2}$) 706.0687 10000.0 0%
Training Set (Avg.) 18.8768 100.0%
Table 2: Long-horizon prediction metrics on the DTDB dataset.
To quantify long-term predictive performance, we use the following metrics: i) NIQE quality score (Mittal et al., 2012)): A common image quality metric assessing the meaningfulness of LDS predictions. Note that some poor predictions (e.g., when most pixel values blow up or converge to 0) can produce unbounded NIQE scores, so we substitute them with 10000; ii) Moving Ratio  [^1] Moving is defined as $F_{M}-F_{M-2}||_{1}>T$, where $\{F_{i}\}_{i=1}^{M}$ is a sequence of $M$ image frames and $T=9$.: The ratio of sequences whose final frames still exhibit the dynamic pattern, indicating whether LDS evolution preserves it.
We report the above metrics for the 285 sequences from the DTDB dataset in Table. 2. The results show that SC best preserves long-term dynamic patterns while maintaining high prediction accuracy, particularly at 500-step predictions, which are nearly comparable to the training set images. This is likely since SC can learn marginally-stable systems ($\varepsilon=0$) or stable systems with slow energy decay ($\varepsilon\rightarrow 0^{+}$), while LS tends to produce unstable long-horizon predictions, and SOC is bound to learn strictly stable systems that might be overly conservative and produce vanishing values. Refer to sample frame predictions in Fig. 8 of Appendix C. We also provide additional experiments and insights into SC’s stability-expressivity tradeoff introduced by the choice of $\varepsilon$ in Appendix C.
@@ 5.5 SC’s benefits extend to nonlinear systems when combined with the Koopman Operator
We evaluated the effectiveness of combining the Koopman Operator theory (Koopman, 1931), SC can be effectively and efficiently applied to highly nonlinear robot manipulation tasks by evaluating SC on DexManip (Rajeswaran et al., 2018b), a robotic dexterous manipulation dataset containing demonstrations of four complex tasks using a 23-DoF multi-fingered hand. Though the original state evolution is nonlinear, it was recently demonstrated that the evolution can be encoded using a linear system when “lifted" to a higher-dimensional space (e.g., $r\approx 750$) that approximates the Koopman invariant subspace Han et al. (2023). Four tasks are briefly described as follows: i) Pen Reorientation: The robot hand is tasked with orienting a pen to a random target rotation angles, ii) Door Opening: The robot hand is tasked with opening a door from a random starting position, iii) Tool Use: The robot hand is tasked with picking up the hammer and driving the nail into the board placed at a random height, and iv) Object Relocation: The robot hand is tasked with moving a cylinder to a random target location (see Appendix. F for the details of each dataset). Similar to the DTDB dataset, only LS, SOC, and SC are applicable due to the high dimensionality of the lifting space.
Computation Time (sec, $\downarrow$) Memory Usage (MB, $\downarrow$) Success Rate (%, $\uparrow$) Safety Rate (%, $\uparrow$   )
@@
We first report the computation time and memory usage in Table. 3. As seen, SC also alleviates the computational and memory burden compared to the SOC method. Note that the values for LS method are marked in gray as this method does not guarantee stability. We then show the average reconstruction error on the DexManip datasets in Fig. 5, where SC demonstrates the best performance and LS results in significantly larger errors on the Tool dataset due to its lack of stability.
Figure 5: Average reconstruction error on the Reorient, Door, Tool, Reloc, Tool_short, and Reloc_with_failures datasets. Solid lines represent mean values, and shaded areas indicate quartile values over the rollout trajectories.
To further underscore the importance of ensuring stability when learning LDSs for robotic dexterous manipulation tasks in the DexManip datasets, we also compute the task success rates and the safety rates (both reported in Table. 3) for each task when tracking the generated robot hand trajectories. The definition of these two metrics are included in Appendix. F.
@@
5.6 SC offers robustness to unsuccessful and shorter demonstrations
In real-world settings, the collected demonstrations may be imperfect, unlike those in the original DexManip datasets. Hence, we introduce two additional datasets: (i) Tool_short, where each demonstration is shorter, posing challenges in learning stable LDSs with a limited horizon, and (ii) Reloc_with_failures, which includes a few failed demonstrations (see Appendix. F for more details).
Computation Time (sec, $\downarrow$) Memory Usage (MB, $\downarrow$) Success Rate (%, $\uparrow$) Safety Rate (%, $\uparrow$   )
@@ 5.6 SC offers robustness to unsuccessful and shorter demonstrations
Table 4: SC can efficiently learn stable manipulation skills from shorter and failed demonstrations.
@@
From Table. 4, and it is evident that SC is more efficient than SOC in terms of compute and memory. In Tool_short and Reloc_with_failures of Fig. 5, we show the average reconstruction error over the original uncorrupted datasets. From these results, we can observe that the LDSs learned by the LS approach tend to be highly unstable, as evidenced by the significantly increasing error over time. A similar conclusion can be drawn from the task success rates and safety rates (both reported in Table. 4), where SC demonstrates significantly superior performance. This indicates the importance of enforcing stability in learned LDSs to effectively handle corrupted demonstrations.
@@
We also reveal the connections between SC and the Koopman Operator’s spectral properties. Following the algorithm described in Rowley et al. (2009), we can rollout the systems with only subsets of the Koopman modes and visualize the hand motions to investigate the effectiveness of SC upon each subset. See detailed experiments in Appendix. D, where we demonstrated how SC effectively stabilizes unstable modes and keeps stable modes.
6 Conclusions
Spectral clipping (SC) is a simple yet surprisingly-effective post-hoc method for ensuring the stability of learned linear time-invariant dynamical systems through extensive experiments across multiple domains. Further, SC can be combined with Koopman-based methods to learn nonlinear dynamical systems and ensure open-loop predictive stability. Key benefits of SC include i) significant gains in computation efficiency without sacrificing prediction accuracy, ii) stable long-term predictions, and iii) robustness to datasets with failed or shorter demonstrations. A key limitation of this work is that we limit our investigation to deterministic systems. Extending these experiments to stochastic systems and automating the choice of the stability margin $\varepsilon$ are promising avenues for future work.